OpenStudy (anonymous):

cost, revenue, and profit A company produces a toy for which the variable cost is $12.30 per unit and the fixed costs are $98,000. The toy sells for $17.98. Let x be the number of units produced and sold. (a) The total cost for a business is the sm of the variable cost and the fixed costs. Write the total cost for C as a function of the number of units produced. (B) Write the revenue R as a function of the number of units sold. (C) Write the profit P as a function of the number of units sold. (Note: P = R - C).

OpenStudy (anonymous):

It kinda spells it out for you how to do it. What part are you stuck on?

OpenStudy (anonymous):

For a) total cost = variable cost + fixed cost. variable cost = $12.30x, etc.

OpenStudy (anonymous):

@emb27 start with: a) total cost = variable cost + fixed cost C = 12.30x + ...

OpenStudy (anonymous):

Then how do I write the function for B & C?

OpenStudy (anonymous):

b) Write the revenue R as a function of the number of units sold. R = 17.98x

OpenStudy (anonymous):

Can you finish the last one?

OpenStudy (anonymous):

Known the important concept of Profit: P = R - C = .... ?

OpenStudy (anonymous):

P = 17.98(12.30) $221.15

OpenStudy (anonymous):

Did you read the above posts?

OpenStudy (anonymous):

Yes. I answered what I thought profit was. P = 17.98(12.30) = $221.15. Correct?

OpenStudy (anonymous):

What kind of operation is it? P = R - C

OpenStudy (anonymous):

What's R? What's C?